OpenStudy (anonymous):

GIVE MEDAL!! In right ∆ABC, m B = 90°, m C = 40°, and BC = 10. What are the other two side lengths of the triangle? AC = 13.1, AB = 8.4 AC = 18.7, AB = 15.8 AC = 14.3, AB = 10.2 AC = 15.6, AB = 11.9

OpenStudy (igreen):

We can use Sine, Cosine, and Tangent to find the remaining sides..

OpenStudy (igreen):



OpenStudy (igreen):

\(\sf Sine=\dfrac{Opposite}{Hypotenuse}\) \(\sf Cosine=\dfrac{Adjacent}{Hypotenuse}\) \(\sf Tangent = \dfrac{Opposite}{Adjacent}\) Here, we are given the adjacent side, and we want to find the opposite and the hypotenuse.

OpenStudy (igreen):

Hehe, it would be if you got it yourself :P

OpenStudy (anonymous):

o..k

OpenStudy (anonymous):

umm..

OpenStudy (igreen):



OpenStudy (igreen):

So to find the hypotenuse, we can use cosine. \(\sf Cosine=\dfrac{Adjacent}{Hypotenuse}\) Plug in what we know: \(\sf cos(90^o) = \dfrac{10}{x}\)

OpenStudy (igreen):

Hmm..I think that's right..I'm not sure about which angle to use..I haven't done this in a while..

OpenStudy (anonymous):

its ok

OpenStudy (igreen):

Okay, 90 degrees is wrong..

OpenStudy (igreen):

Hold on a minute

OpenStudy (anonymous):

k

OpenStudy (igreen):

Okay, I think we use 40 degrees..

OpenStudy (igreen):

So: \(\sf cos(40^o) = \dfrac{10}{x}\) Type in \(\sf cos(40^o)\) into your calculator, tell me what you get.

OpenStudy (anonymous):

i dont have a scientific one ..its ok ill guess but thx for help ill give tht medal :)

OpenStudy (igreen):

You can use google, just type in "cos 40 degrees"

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

a big decimal so ill put 0.766

OpenStudy (igreen):

Yep, so we have: \(\sf 0.766 = \dfrac{10}{x}\) Solve for 'x'

OpenStudy (anonymous):

so i times the decimal and 10?

OpenStudy (anonymous):

or divide

OpenStudy (igreen):

Well, you can switch 'x' and 0.766 around. \(\sf x = \dfrac{10}{0.766}\) So just divide

OpenStudy (anonymous):

13.1?

OpenStudy (anonymous):

it was a big decimal so i round it and got tht

OpenStudy (igreen):

Sorry, my internet just went out for a few seconds there.

OpenStudy (igreen):

Yep, 13.1 is corret.

OpenStudy (igreen):

correct*

OpenStudy (igreen):

That's our hypotenuse.

OpenStudy (anonymous):

its a!!!

OpenStudy (igreen):

Yes, I think that's correct.

OpenStudy (anonymous):

thx for the help :)

OpenStudy (igreen):

Np